40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 25 19:55:59 2009, Jenne, Update, Adaptive Filtering, Subtraction of the microseism using Adaptive Filtering! OAF_running_WORKING_25Sept2009.png
    Reply  Mon Sep 28 02:24:05 2009, rana, Update, Locking, MC1/3 Dewhitening found OFF: Turned back ON C1SUS_SRM_XYCOM1.pngUntitled.png
Message ID: 2011     Entry time: Mon Sep 28 02:24:05 2009     In reply to: 2004
Author: rana 
Type: Update 
Category: Locking 
Subject: MC1/3 Dewhitening found OFF: Turned back ON 

While trying to make the OAF work, I found that the XYCOM switches for MC1/3 have been set in the bad way for awhile. This means that the hardware filters were bypassed and that MC1 & MC3 were moving around too much at high frequency and possibly causing trouble with the locking. I have put them back into the default position.

On Friday, Jenne and I were playing around with turning the dewhitening off/on to see if it efffected the OAF stability. At the time, I didn't pay too much attention to what the state was. Looks like it was in the wrong state (hardware bypassed) when we found it. For the OAF work, we generally want it in that bypassed state, but its bad because it makes noise in the interferometer. The bits in question are bits 16-23 on the XYCOM screen.

I have updated the snapshot and set the screen in the appropriate settings. I used a swept sine measurement to verify the filter state. In the attached plot, green corresponds to XYCOM green and red corresponds to red.

Attachment 1: C1SUS_SRM_XYCOM1.png  39 kB  | Hide | Hide all | Show all
C1SUS_SRM_XYCOM1.png
Attachment 2: Untitled.png  37 kB  | Show | Hide all | Show all
ELOG V3.1.3-